# Optimal, fast, and robust inference of reionization-era cosmology with the 21cmPIE-INN

### v2

@@ Abstract
Modern machine learning will allow for simulation-based inference from reionization-era 21cm observations at the Square Kilometre Array. Our framework combines a convolutional summary network and a conditional invertible network through a physics-inspired latent representation. It allows for an optimal and extremely fast determination of the posteriors of astrophysical and cosmological parameters. The sensitivity to non-Gaussian information makes our method a promising alternative to the established power spectra.
@@ 2.5 Validation
This equation allows us to check for self-consistent sampling. If one samples from the correct posterior, Eq.(11) holds for any form of the posterior. Any violation of this equality indicates a problem in the sampling. If there is no violation then one can claim that the obtained posteriors are optimal and without bias. The authors of Ref. [70] propose the SBC algorithm
@@ 3.2 Mock data
The same bottom line can also be extracted from the simulation-based calibration in Fig. 6. We show histograms for a sample size of 10000 and 15 bins. Self-consistent sampling leads to a uniform distribution. The shaded region is the expected variation based on the 99% quantile of a Binomial. Common failure modes are overestimation, leading to a $\cup$-shape, and underestimation, leading to a $\cap$-shape. An asymmetry points towards a systematic bias. In our case, the histograms for the pure simulation show the expected variance, implying self-consistent sampling. After adding noise, all parameters except for $E_{0}$, and to a much smaller degree $\zeta$, remain well-calibrated. For $E_{0}$ the $\cup$-shape, implies under-confidence, corresponding to the issues observed in Fig. 5. Unlike standard MCMC analysis, where SBC is computationally not feasible, we can claim that our posteriors have the correct coverage and no bias.
@@ 4 Outlook
The summary vector that links the two networks of the 21cmPIE-INN is initialized to the parameters of interest, but adapted by the joint training with the cINN to guarantee an optimal inference. This optimization, the impact of the form and size of the summary vector, its stability, and its benefits in terms of explainable AI is an interesting avenue for further investigation. Similarly, our method allows for training on augmented datasets, for example with different experimental performance, opening further possibilities to analyze realistic SKA data.
Ultimately, the proposed approach facilitates rapid and straightforward simulation-based inference of likelihood constraints for cosmology. The generation of full posteriors from a given 21cm light cone takes a few seconds, most of this time used for reading in the complex data. This speed, combined with the ability to capture non-Gaussian information optimally, distinguishes this method from alternative inference approaches.

### v3

@@ Abstract
Modern machine learning will allow for simulation-based inference from reionization-era 21cm observations at the Square Kilometre Array. Our framework combines a convolutional summary network and a conditional invertible network through a physics-inspired latent representation. It allows for an efficient and extremely fast determination of the posteriors of astrophysical and cosmological parameters, jointly with well-calibrated and on average unbiased summaries. The sensitivity to non-Gaussian information makes our method a promising alternative to the established power spectra.
@@ 2.5 Validation
This equation allows us to check for self-consistent sampling. If one samples from the correct posterior, Eq.(11) holds for any form of the posterior. Any violation of this equality indicates a problem in the sampling. If there is no violation then one can claim that the posterior is on average without bias and uses all the information from the summary. However, different biases in certain regions may cancel each other and the summary might not be optimal. The authors of Ref. [70] propose the SBC algorithm
@@ 3.2 Mock data
The same bottom line can also be extracted from the simulation-based calibration in Fig. 6. We show histograms for a sample size of 10000 and 15 bins. Self-consistent sampling leads to a uniform distribution. The shaded region is the expected variation based on the 99% quantile of a Binomial. Common failure modes are overestimation, leading to a $\cup$-shape, and underestimation, leading to a $\cap$-shape. An asymmetry points towards a systematic bias. In our case, the histograms for the pure simulation show the expected variance, implying self-consistent sampling. After adding noise, all parameters except for $E_{0}$, and to a much smaller degree $\zeta$, remain well-calibrated. For $E_{0}$ the $\cup$-shape, implies under-confidence, corresponding to the issues observed in Fig. 5. Unlike standard MCMC analysis, where SBC is computationally not feasible, we can claim that our posteriors have the correct coverage and on average no bias.
@@ 4 Outlook
The summary vector that links the two networks of the 21cmPIE-INN is initialized to the parameters of interest, but adapted by the joint training with the cINN to guarantee an optimal inference, for the given underlying network. This optimization, the impact of the form and size of the summary vector, its stability, and its benefits in terms of explainable AI is an interesting avenue for further investigation. Similarly, our method allows for training on augmented datasets, for example with different experimental performance, opening further possibilities to analyze realistic SKA data.
Ultimately, the proposed approach facilitates rapid and straightforward simulation-based inference of likelihood constraints for cosmology. The generation of full posteriors from a given 21cm light cone takes a few seconds, most of this time used for reading in the complex data. This speed, combined with the ability to capture non-Gaussian information and learn optimal summaries (meaning well-calibrated with SBC and on average unbiased), distinguishes this method from alternative inference approaches.